OpenStudy (anonymous):

find the intersections: x squared plus y squared = 1 y=x

OpenStudy (anonymous):

In that case, 2x^2 = 1 x^2 = 1/2 x = +/- 1/SQRT(2)

OpenStudy (anonymous):

thanks bro, that was def way more of an advanced answer than i thought it would be